OpenStudy (madmerc):

PLEASE HELP!!! Write the equation of the line that passes through (−3, 5) and (2, 10) in slope-intercept form.

Mehek (mehek14):

slope formula \(\tt{\dfrac{y_2-y_1}{x_2-x_1}}\)

Mehek (mehek14):

\(\tt{x_1=-3\\y_1=5\\x_2=2\\y_2=10}\)

OpenStudy (madmerc):

The choices are: y = x + 8 y = x − 8 y = −5x − 10 y = −5x + 20

Mehek (mehek14):

ok so let's find the slope \(\tt{\dfrac{10-5}{2-(-3)}=\dfrac{5}{5}=1}\)

Mehek (mehek14):

so slope is x

Mehek (mehek14):

\(\tt{y=x+b}\) we need to find the y-intercept

Mehek (mehek14):

we can use any of the ordered pairs.

OpenStudy (madmerc):

so (-3, 5) would be 5 = -3 + b ?

Mehek (mehek14):

yes so solve for b

OpenStudy (madmerc):

is it b = 8 ?

Mehek (mehek14):

yup

Mehek (mehek14):

so slope is 1x or just x y-intercept is 8

Mehek (mehek14):

that makes the equation \(\tt{y=x+8}\)

OpenStudy (madmerc):

oh, okay! thank you

Mehek (mehek14):

yw :]